how do you solve
x2-3x+2?
8x(Squared)-4x=24?
2x(squared)+9x+4?
16x(squared)-8x=-1

Answer by checkley77(12844)

x^2-3x+2?
(x-2)(x-1)
8x(Squared)-4x=24?
8x^2-4x-24=0
4(2x^2-x-6)=0
4(2x+3)(x-2)=0
2x+3=0
x=3/2 ans.
x-2=0
x=2 ans.
2x(squared)+9x+4?
2x^2+9x+4
(2x+1)(x+4)
16x(squared)-8x=-1
16x^2-8x+1=0
(4x-1)(4x-1)=0
4x-1=0
4x=1
x=1/4 ans.